OpenStudy (anonymous):

help please!! This is a spinner used in a board game. What is the probability that the spinner will land on a multiple of 3 and 4? A. 1/4 B. 1/2 C. 1/8 D. 3/8

OpenStudy (anonymous):

@jim_thompson5910 can you help me with this one

OpenStudy (anonymous):

What numbers are on the spinner?

OpenStudy (anonymous):

sorry ! forgot about the picture

jimthompson5910 (jim_thompson5910):

how many numbers on that spinner are multiples of 3 and 4?

OpenStudy (anonymous):

only 12

jimthompson5910 (jim_thompson5910):

12, what else

OpenStudy (anonymous):

all of them but 10

OpenStudy (anonymous):

and 22

jimthompson5910 (jim_thompson5910):

no, not multiples of 3 or 4 multiples of 3 AND 4

OpenStudy (anonymous):

thats what i thought , thats why i said 12

jimthompson5910 (jim_thompson5910):

there are others though, not just 12

OpenStudy (anonymous):

i dont understand , sorry.

jimthompson5910 (jim_thompson5910):

wouldn't 24 be another one? there's one more

OpenStudy (anonymous):

54?

OpenStudy (anonymous):

or 64 im not sure which

OpenStudy (anonymous):

its neither! SORRY 36

jimthompson5910 (jim_thompson5910):

54 doesn't work (4 isn't a factor) 64 doesn't work (3 isn't a factor)

jimthompson5910 (jim_thompson5910):

yep 36

jimthompson5910 (jim_thompson5910):

so there are 3 that work (out of 8), giving you 3/8

OpenStudy (anonymous):

thank you , sorry i didnt get what you was saying at first but i multiplied 3*4=12+3*4=24+3*4=36 got it now :)

jimthompson5910 (jim_thompson5910):

ok that's great